OpenStudy (anonymous):

WRITE THE SLOPE INTERCEPT EQUATION FOR THE LINE THAT PASSES THROUGH (-4,6)AND (2,-6) PLEASE SHOW ALL WORK

OpenStudy (anonymous):

I DID -6-6/2-4=0/-2

OpenStudy (amistre64):

-6-6 not= 0

OpenStudy (amistre64):

its easier to line them up vertically to me: (-4, 6 ) - ( 2 ,-6) ----------- -6, 12 then stack y/x: 12/-6

OpenStudy (amistre64):

then we can get to the slope intercept using either point as a base; along with the slope we just found y = mx -mPx + Py y = -2x --2(2)+-6 y = -2x +2(2)-6 y = -2x +4-6 y = -2x -2

OpenStudy (anonymous):

I GOT IT THANKS

OpenStudy (amistre64):

good job :)

OpenStudy (anonymous):

BUT THE FORMULA IS Y2-Y1/X2-X1

OpenStudy (anonymous):

I C NEVER MIND